Without using finance calculator,
(a) What sum of money will become RM 15,000 in three years at a simple interest rate of 8% per annum?
(b) A certain sum of money is invested now. This investment will be worth RM5500 after fifteen months and RM5800 after twenty-four months. Find the original principal and the simple interest rate that was offered.
(c) A finance company charges a simple interest of 18% per year for short-term loans. If a loan of RM 15,000 was charged RM 337.50, how long was the term of the loan?
1
Expert's answer
2020-05-27T10:13:13-0400

a) If the initial investment is x, after three years it will become:

"x*(1+0.08*3)=15000"


Which means that:

"x=\frac{15000}{1+0.08*3}=\frac{15000}{1.24}\approx12096.77"


b) If the initial investment is x and interest rate is r, investment after fifteen and twenty four month will be equial to:

"x*(1+\frac{15}{12}*r)=5500\\\\x*(1+\frac{24}{12}*r)=5800"

Which means that original principal is RM 5 000, offered simple interest rate is 8%.


c) If term of the loan is t years, charged interest will be:

"15000*0.18*t=337.50"

Which means that term of the loan is:

"t=\frac{337.5}{15000*0.18}=0.125" years or 1.5 month.